A model car of mass 4kg moves in a straight line along horizontal ground. The resistance to motion of the car has magnitude 4N. The car batteries provide a constant power of 20watts. Calculate:
(i) the acceleration of the car at an instant when it's speed is 2m/s.
(I) when the car passes point A, it is moving with constant spread Vm/s. Show that V=5m/s
1
Expert's answer
2019-10-24T09:36:42-0400

Mass of car"=4 Kg"

Power of batteries"=20 W"

Resistive force"=4N"

"(i)" Velocity of car"=2 m\/sec"

Resistive power"=Fv=4\times 2=8 W"

Net power"=20W-8W=12W"

"P_{net}=F_{net}v\implies F_{net}=\frac{12}{2}=6N"

"F_{net}=ma\implies a=\frac{F_{net}}{m}=\frac{6}{4}=1.5 m/sec^2"

"(l)" When speed is constant ,Resistive power"=" Battery power

So,"P(" batteries)"=P" (resistance)"=P=20 N" ;

"P=Fv\implies 20=4\times v\implies v=\frac{20}{4}=5m/sec"